# Superconvergent and Divergence-Free Finite Element Methods for Stokes Equation

### v1

@@ 5. Numerical examples
$2^{-1}$ $(0,0)$ 6.762e-04 9.787e-03 2.942e-01
$2^{-2}$ $(0,0)$ 3.110e-04 1.12 4.960e-03 1.06 1.649e-01 0.84
$2^{-3}$ $(0,0)$ 1.399e-04 1.15 2.431e-03 0.95 8.501e-02 0.96
$2^{-4}$ $(0,0)$ 6.719e-05 1.06 1.260e-03 0.95 4.284e-02 0.99
$2^{-5}$ $(0,0)$ 3.320e-05 1.02 6.397e-04 0.98 2.146e-02 1.00
$2^{-1}$ $(1,0)$ 2.076e-04 2.431e-03 2.942e-01
$2^{-2}$ $(1,0)$ 6.700e-05 1.63 5.715e-04 2.09 1.649e-01 0.84
$2^{-3}$ $(1,0)$ 1.790e-05 1.90 1.541e-04 1.89 8.501e-02 0.96
$2^{-4}$ $(1,0)$ 4.551e-06 1.98 3.975e-05 1.96 4.284e-02 0.99
Table 4. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
$2^{-1}$ $(0,0)$ 4.167e-04 4.576e-03
$2^{-2}$ $(0,0)$ 1.565e-04 1.41 2.880e-03 0.67
$2^{-3}$ $(0,0)$ 4.400e-05 1.83 1.539e-03 0.90
$2^{-4}$ $(0,0)$ 1.143e-05 1.95 7.836e-04 0.97
$2^{-5}$ $(0,0)$ 2.889e-06 1.98 3.937e-04 0.99
$2^{-1}$ $(1,0)$ 9.624e-05 1.894e-03
$2^{-2}$ $(1,0)$ 1.493e-05 2.69 5.478e-04 1.79
$2^{-3}$ $(1,0)$ 2.035e-06 2.88 1.486e-04 1.88
$2^{-4}$ $(1,0)$ 2.627e-07 2.95 3.808e-05 1.96

### v2

@@ 5.2. Singular problem with slip boundary conditions
$1/14$ $(0,0)$ 1.866e-01 1.456e+00 1.364e+00
$1/28$ $(0,0)$ 1.156e-01 0.70 1.023e+00 0.51 1.006e+00 0.44
$1/56$ $(0,0)$ 7.488e-02 0.63 7.737e-01 0.40 7.250e-01 0.47
$1/112$ $(0,0)$ 4.874e-02 0.62 6.093e-01 0.35 5.170e-01 0.49
$1/224$ $(0,0)$ 3.153e-02 0.63 4.865e-01 0.33 3.693e-01 0.49
$1/14$ $(1,0)$ 7.867e-02 1.791e+00 8.350e-01
$1/28$ $(1,0)$ 5.253e-02 0.59 1.396e+00 0.36 6.071e-01 0.46
$1/56$ $(1,0)$ 3.455e-02 0.60 1.097e+00 0.35 4.423e-01 0.46
$1/112$ $(1,0)$ 2.242e-02 0.63 8.638e-01 0.34 3.257e-01 0.44
$1/224$ $(1,0)$ 1.440e-02 0.64 6.820e-01 0.34 2.434e-01 0.42
The numerical results for Example 5.3 are reported in Table 4. The corner singularity implies that $\boldsymbol{\sigma},p\in H^{1/3-\varepsilon}(\Omega)$. Hence, on uniform meshes, one expects
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}=\mathcal{O}(h^{1/3}),\quad\|p-p_{h}\|=\mathcal{O}(h^{1/3}),\quad\|\boldsymbol{u}-\boldsymbol{u}_{h}\|=\mathcal{O}(h^{2/3}).$
The observed rates are consistent with these predictions. Fig. 2 shows the velocity field on the L-shaped domain for two numerical solutions and the exact solution. In particular, the effect of the re-entrant corner is clearly visible. Although the convergence order of $\boldsymbol{u}_{h}$ is limited by the regularity, BDM1 with $(k,\ell)=(1,0)$ still performs better than RT0 with $(k,\ell)=(0,0)$ away from the singularity.
6. Conclusions and Future Work
In this paper, a family of pointwise divergence-free mixed finite element methods is developed for the Stokes equation. By using a distributional discretization of the vector Laplacian, the method gives a consistent and stable scheme even though the velocity space is only $H(\operatorname{div})$-conforming. The central mechanism is the weak deviatoric gradient operator $\operatorname{dev}{\rm grad\,}_{w}$, which uses tangential–normal continuous traceless tensor elements to preserve discrete integration-by-parts identities and commuting properties. In addition, a supercloseness result is established for the velocity and pressure approximations, and a post-processed discrete velocity achieves super-convergent rate.
The framework also points to several directions for future work:
  1. (1)
    Brinkman and Navier–Stokes equations: The $\operatorname{dev}{\rm grad\,}_{w}$ framework is well suited to the Brinkman model, where the weak deviatoric gradient operator can naturally handle the transition between Darcy and Stokes regimes. For high-Reynolds-number Navier–Stokes flows, additional stabilization is still needed when convection dominates.
  2. (2)
    Linear elasticity: The link between stress-like variables and divergence constraints suggests extensions to mixed elasticity. Symmetric tensor spaces in this framework may lead to robust elements for nearly incompressible materials.
  3. (3)
    Boundary conditions and regularity: The weak operators can be adapted to slip and traction boundary conditions through modified boundary trace terms. It is also important to study superconvergence and adaptive finite element methods on nonconvex domains.